View Issue Details Jump to Notes ] Issue History ] Print ]
IDProjectCategoryView StatusDate SubmittedLast Update
0000337Hall D OnlineMagnetspublic2013-04-03 16:242013-07-18 09:14
Assigned Towolin 
PlatformOSOS Version
Summary0000337: Refurbish Danfysik power supply for production running
DescriptionRefurbish our boards, return Hall C boards, replace other components as needed, get Danfysik techs out here to check everything.
TagsNo tags attached.
Work by outside group for Hall DNo
Estimated man-weeks2
Percentage complete0
Actual man-weeks0
Attached Files

- Relationships

-  Notes
wolin (administrator)
2013-04-03 16:25

Danfysik techs will be here this summer to work on supplies in other halls. Tim is organizing for them to work on our supply as well.
wolin (administrator)
2013-05-16 17:10

Need to tune for 25H instead of 2H (done for Hall A). Need to program ramp profile to get smooth voltage profile, including no overshoot. Probably need timer to inhibit quench detection when the contacter is first closed for (say) 100 ms.
wolin (administrator)
2013-06-14 10:19
edited on: 2013-06-14 10:20

Parts for tuning for 25H on the way from Danfysik. Giles figured out how to program hardware ramp profiles. Dave B will think about doing the same in the PLC.

Note that the hardware has many constraints, including the number of points per profile and the number of profiles stored. The PLC has no constraints. Note that we'll need profiles for ramping up AND down. Further the PLC could generate ramp profiles on the fly based on start/end currents and number of steps, so nothing would need to be stored.

wolin (administrator)
2013-06-14 10:21

Waiting for parts to arrive.
wolin (administrator)
2013-06-25 16:59

Parts here, not sure if we'll install them before or after solenoid test.
wolin (administrator)
2013-07-18 09:13

Since we had innumerable delays in the cryo plant all parts were installed, FPGA programs were updated, etc.
wolin (administrator)
2013-07-18 09:14

We still probably want Danfysik techs to come out and go over the whole power supply.

- Issue History
Date Modified Username Field Change
2013-04-03 16:24 wolin New Issue
2013-04-03 16:24 wolin Work by outside group for Hall D => No
2013-04-03 16:24 wolin Estimated man-weeks => 2
2013-04-03 16:24 wolin Percentage complete => 0
2013-04-03 16:24 wolin Actual man-weeks => 0
2013-04-03 16:24 wolin Status new => assigned
2013-04-03 16:24 wolin Assigned To => wolin
2013-04-03 16:25 wolin Note Added: 0000508
2013-04-03 16:25 wolin Status assigned => acknowledged
2013-05-16 17:10 wolin Note Added: 0000517
2013-06-14 10:19 wolin Note Added: 0000533
2013-06-14 10:20 wolin Note Edited: 0000533
2013-06-14 10:21 wolin Note Added: 0000534
2013-06-14 10:21 wolin Status acknowledged => feedback
2013-06-25 16:59 wolin Note Added: 0000545
2013-07-18 09:13 wolin Note Added: 0000553
2013-07-18 09:14 wolin Note Added: 0000554

Copyright © 2000 - 2020 MantisBT Team
Powered by Mantis Bugtracker